OpenStudy (anonymous):

What is the mole fraction of solute and the molal concentration for an aqueous solution that is 16.0% NaOH by mass?

OpenStudy (xishem):

Assume 100g of solution... \[16.0g\ NaOH*\frac{1mol\ NaOH}{40.00g\ NaOH}=0.400mol\ NaOH\]\[84.0g\ H_2O*\frac{1mol\ H_2O}{18.02g\ H_2O}=4.66mol\ H_2O\] The molefraction is the ratio of amount of constituent to total amount of substance. So...\[0.400 : (0.400+4.66)=0.400:5.06=0.0791\]And the molal concentration is defined as...\[b_i=\frac{n_i}{m_{solvent}}\]Where... b_i is the molal concentration of in in the solution n_i is the number of mols of i in the solution m_solvent is the mass of the solvent in kilograms \[b_i=\frac{0.400mol\ NaOH}{0.084kg\ H_2O}=4.76m\]